OpenStudy (anonymous):

You need to determine where to place the beams so that the chains are fastened to the rollercoaster at a height of 25 feet. 1. 3. Write the equation you would need to solve to find the horizontal distance each beam is from the origin

OpenStudy (anonymous):

Needs more information.

OpenStudy (anonymous):

I can give more information

OpenStudy (anonymous):

What is to say the answer isn't 4? Basically, there are other facts here that need to be given for there to be a single answer.

OpenStudy (anonymous):

This whole sheet confuses me. not just that question.

OpenStudy (anonymous):

\[ x^2+y^2=r^2 \]So\[ x^2+25^2=30^2 \]

OpenStudy (anonymous):

\[ x = \sqrt{30^2-25^2} \]

OpenStudy (anonymous):

So, do they cancel and it equals 16.6?